# Incentive-Compatible Vertiport Reservation in Advanced Air Mobility: An Auction-Based Approach

### v2

@@ Abstract
The rise of advanced air mobility (AAM) is expected to become a multibillion-dollar industry in the near future. Market-based mechanisms are touted to be an integral part of AAM operations, which comprise heterogeneous operators with private valuations. In this work, we study the problem of designing a mechanism to coordinate the movement of electric vertical take-off and landing (eVTOL) aircraft, operated by multiple operators each having heterogeneous valuations associated with their fleet, between vertiports, while enforcing the arrival, departure, and parking constraints at vertiports. Particularly, we propose an incentive-compatible and individually rational vertiport reservation mechanism that maximizes a social welfare metric, which encapsulates the objective of maximizing the overall valuations of all operators while minimizing the congestion at vertiports. Additionally, we improve the computational tractability of designing the reservation mechanism by proposing a mixed binary linear programming approach that is based on constructing network flow graph corresponding to the underlying problem.
@@ 1 Introduction
Despite the widespread optimism surrounding AAM, the design of regulatory policies remains an open problem. While ideas from conventional air traffic framework could be leveraged (Bertsimas and Patterson, 1998, 2000; Bertsimas et al., 2011; Roy and Tomlin, 2007; Odoni, 1987), they often fall short in accommodating the dynamic and adaptable nature of AAM operations (Bichler et al., 2023). Unlike traditional aviation, AAM requires a more flexible regulatory approach to be able to handle on-demand operations comprising of broader spectrum of arrival and departure points (Skorup, 2019; Seuken et al., 2022).
@@
Administrative management methods prevalent in conventional airspace, such as grand-fathering rights, flow management, and first-come-first-serve, prove ineffective for AAM operations (Guerreiro et al., ; Evans et al., ). These approaches fail to elicit the heterogeneous private valuations (arising from different aircraft specifications, demand realization, etc.) different operators have on using AAM resources. Furthermore, they risk fostering inefficient and anti-competitive outcomes, as evidenced in traditional airspace operations (Dixit et al., 2023). Recognizing the need for tailored regulation, the Federal Aviation Administration (FAA) is actively developing a clean-slate congestion management framework for AAM operations to ensure efficiency, fairness, and safety (Administration, 2023).
@@
Market-based congestion management mechanisms have been proposed as potential solutions for AAM operations (Chin et al., 2023b; Qin and Balakrishnan, 2022; Wang et al., 2023; Evans et al., ; Skorup, 2019; Seuken et al., 2022). In this paper, we study an auction-based mechanism for AAM congestion management.
@@
Figure 1: Schematic representation of the air traffic network with multiple service providers, each operating in non-overlapping operation regime, tasked with coordinating the movement of aircraft of various fleet operators between vertiports in its domain. Each vertiport has a constraint on the number of arriving aircraft, departing aircraft, and parked aircraft.
Even in conventional airspace management, auction-based mechanisms are extensively studied such as (Ball et al., 2018; Basso and Zhang, 2010; Carlin and Park, 1970; Mehta and Vazirani, 2020), where both theoretical and empirical evidence show their precedence over administrative approaches (Dixit et al., 2023). Despite being a plausible direction, auctioning airspace is rarely adopted in conventional airspace due to intricate and inflexible government regulations in conventional airspace operations (Bichler et al., 2023).
The design of market-based mechanisms that guarantee safety, efficiency, and fairness under the heterogeneous and on-demand nature of AAM operations has remained elusive. Particularly, there are two challenges. First, there are computational challenges associated with computing the auction mechanism. Second, there are challenges to truthfully elicit private information (aka incentive compatibility) from AAM operators. Existing approaches in AAM operations concentrate heavily on tactical deconfliction (Kleinbekman et al., 2018; Bertram and Wei, ), while strategic trajectories are determined through flow management frameworks or other protocols which do not account for economic incentives (Chin et al., 2021; Sun et al., 2023; Chin et al., 2023a, b; Wang et al., 2023; Qin and Balakrishnan, 2022; Guerreiro et al., ; Evans et al., ; Yu et al., 2023). Facing these issues, we propose a vertiport reservation mechanism, based on auction theory, that can be integrated into the airborne automation workflow proposed in (Wei et al., 2023). Particularly, the main motivating question in this work is as follows.
How do we design an incentive-compatible vertiport reservation mechanism that can accommodate the heterogeneous nature of AAM operations, while having good computational performance?
A schematic representation of vertiport reservation problem is presented in Figure 1. Typically, the air transportation network (ATN) for eVTOL aircraft is envisioned to be divided into several contiguous regions, each operated by a service provider (SP) (Skorup, 2019). In this work, we focus on one region of ATN which is managed by a SP, who is responsible for ensuring the efficient, safe, and fair movement of aircraft between various vertiports. The aircraft are owned by fleet operators (FOs), and each FO may own multiple aircraft. Each FO submits a menu of routes for each of its aircraft to SP, for each of which it has private valuation. The SP is then responsible for allocating different aircraft to routes.
The goal of the SP is to maximize a metric of social welfare that is comprised of two objectives: $(i)$ maximizing the overall (weighted) valuations of all FOs [^1] We allow the SP to weigh FOs differently in order to encourage new-comers in this emerging market., and $(ii)$ minimizing excessive congestion at vertiports. Additionally, the SP must $(iii)$ ensure safety by enforcing arrival, departure, and parking capacity constraints at vertiports, and $(iv)$ elicit truthful valuations from heterogeneous FOs in the form of bids. We emphasize that this problem is an “exchange problem”, where some of the resources desired by any FO could be occupied by aircraft of FOs, and a feasible allocation in this setting needs to exchange the resources between FOs while respecting capacity constraints. In contrast, the standard slot allocation problems studied in conventional air traffic literature (cf. (Dixit et al., 2023; Mehta and Vazirani, 2020; Ball et al., 2018; Rassenti et al., 1982; Pertuiset and Santos, 2014; Ball et al., 2020; Bichler et al., 2023)) are “assignment problems” where the slots need to be assigned to airlines and not exchanged between airlines.
@@
In this work, we propose an auction mechanism for the SP that satisfies $(i)-(iv)$. In our proposed mechanism, using the bids submitted by FOs, the SP allocates the resources by maximizing social welfare, subject to constraints. Additionally, we propose a payment mechanism, inspired by the generalized Vickrey–Clarke–Groves (VCG) mechanism (Nisan et al., 2007), that ensures that bids reveal true valuations of FOs. Specifically, each FO is charged a payment based on the externality imposed by them. Externality imposed by any FO is assessed by the difference in the optimal social welfare of remaining FOs when this FO is present versus when it is excluded from the auction environment. We theoretically study properties of the proposed mechanism in terms of incentive compatibility, individual rationality, and social welfare maximization (cf. Theorem 3.3).
@@
There are two computational challenges associated with designing the auction mechanism. First, naively optimizing social welfare over the set of feasible allocations could be computationally challenging. Therefore, we frame the problem as a mixed integer linear program which is further converted to a mixed binary linear program by constructing a network-flow graph which consists of reduced number of discrete variables. Second, the computation of externality for the payment mechanism requires characterizing the set of feasible allocations when an FO is excluded from the auction environment, which is non-trivial as the underlying resource allocation problem is an exchange problem. Therefore, we introduce the idea of pseudo-bids, where we simply set a bid of 00 to an FO while computing the optimal allocation when this FO is excluded from the auction environment.
@@
Notation: We denote the set of real numbers by $\mathbb{R}$, non-negative real numbers by $\mathbb{R}_{+}$, integers by $\mathbb{Z}$, non-negative integers by $\mathbb{Z}_{+}$, and natural numbers by $\mathbb{N}$. For $N\in\mathbb{N}$, we define $[N]:=\{1,2,...,N\}$. The indicator function is denoted as $\mathbf{1}(\cdot)$, which is 1 when ()($\cdot$)( ⋅ ) is true and 0 otherwise. When indexing a vector $b=(b_{1},b_{2},...,b_{N})$, we follow the standard game-theoretic notation: $b_{-i}:=(b_{1},...,b_{i-1},b_{i+1},...,b_{N})$.
A table containing all major notations used in this article is presented in Appendix A.
@@ 2.1 System Model
2.1.1 High-level description of model:
We examine an air transportation network (ATN) comprising multiple vertiports, catering to electric vertical take-off and landing (eVTOL) fleet operators, each operating multiple eVTOL aircraft. Our focus lies on a strategic deconfliction mechanism, complementing the tactical deconfliction algorithms proposed in prior works [(Kleinbekman et al., 2018; Wu et al., 2022; Bertram and Wei, ; Shao et al., 2021)]. We tackle the challenge of efficiently coordinating the arrival and departure of aircraft on the ATN, over a finite time window, by soliciting valuations from fleet operators in the form of bids [^2] Our proposed solution is engineered to operate recurrently, seamlessly spanning consecutive time intervals., while ensuring all operational constraints are met.
At the onset of each time window, every fleet operator (FO) submits a comprehensive “menu” of ”routes” — a detailed list comprising origin-destination pairs along with departure and arrival times for each aircraft in their fleet. Concurrently, for each route in the menu, they include corresponding bids indicating the amount they are willing to pay should their aircraft be allocated that specific route. Subsequently, the service provider (SP) undertakes the task of computing a feasible allocation and corresponding payments, executing them seamlessly throughout the time window. With allocations finalized, granted aircraft are free to proceed to their desired destinations.
In situations where vertiports face congestion, reaching full parking capacity, any additional arrivals trigger the simultaneous departure of an aircraft from that vertiport. This dynamic transforms the problem into an “exchange problem” rather than the conventional “assignment problem” typically encountered in other air traffic allocation scenarios (Dixit et al., 2023; Mehta and Vazirani, 2020; Ball et al., 2018; Rassenti et al., 1982; Pertuiset and Santos, 2014; Ball et al., 2020; Bichler et al., 2023).
2.1.2 Detailed description of model:
We denote the set of vertiports by $R$, the set of FOs by $F$, and the set of eVTOL aircraft by $A$. We consider the problem of scheduling aircraft to and from vertiports over a time window of $H$ time slots.
@@ 2.1.2 Detailed description of model:
At any time $t\in[H]$, each vertiport $r\in R$ has three kind of capacity constraints: (i) arrival capacity constraints, denoted by $\textsf{arr}(r,t)\in\mathbb{Z}_{+}$, that restrict the number of eVTOLs that can land at vertiport $r$ at time $t$; (ii) departure capacity constraints, denoted by $\textsf{dep}(r,t)\in\mathbb{Z}_{+}$, that restrict the number of eVTOLs that can depart from vertiport $r$ at time $t$; (iii) parking capacity constraints, denoted by $\textsf{park}(r,t)\in\mathbb{Z}_{+}$, that restrict the number of eVTOLs that can park at vertiport $r$ at time $t$.
@@
Let $A_{i}$ be the fleet of aircraft operated by FO $i\in F$, and $A:=\{a_{i,j}|i\in F,j\in A_{i}\}$ be the set of all aircraft using the ATN. Each aircraft $a_{i,j}$ is identified by the following tuple
$a_{i,j}\equiv\left(r^{\textsf{orig}}_{i,j},m_{i,j},\{t^{\textsf{dep}}_{i,j,k},t^{\textsf{arr}}_{i,j,k},v_{i,j,k},b_{i,j,k},r^{\textsf{dest}}_{i,j,k}\}_{k\in m_{i,j}}\right),$ (Aircraft)
where $(i)$ $r^{\textsf{orig}}_{i,j}\in R$ is the origin vertiport of aircraft $a_{i,j}$, $(ii)$ $m_{i,j}$ is the menu of available routes [^3] We include the option to stay parked at the same vertiport in this menu. We shall denote this route by $\varnothing$. The departure time for such an option is set to 00. to aircraft $a_{i,j}$; $(iii)$ any route $k\in m_{i,j}$ implies that aircraft $a_{i,j}$ departs from $r^{\textsf{orig}}_{i,j}\in R$ at time $t^{\textsf{dep}}_{i,j,k}$ to arrive at $r^{\textsf{dest}}_{i,j,k}\in R$ at time $t^{\textsf{arr}}_{i,j,k}$; $(iv)$ $v_{i,j,k}$ denotes the private valuation of aircraft $a_{i,j}$ to choose the route $k\in m_{i,j}$; and $(v)$ $b_{i,j,k}\in\mathbb{R}_{+}$ is the bid submitted by FO $i$ to schedule aircraft $a_{i,j}$ on route $k\in m_{i,j}$. For concise notation, we define $L:=\sum_{i\in F}\sum_{j\in A_{i}}|m_{i,j}|$ to be the number of all possible routes, $L_{i}:=\sum_{j\in A_{i}}|m_{i,j}|$ to be the number of routes available to fleet operator $i\in F$, and $L_{-i}:=\sum_{\ell\in F\backslash\{i\}}\sum_{j\in A_{\ell}}|m_{\ell,j}|$ to be the number of routes available to fleet operators except $i$.
Additionally, for concise notation, we denote the joint bid profile of all aircraft operated by FO $i\in F$ by $B_{i}:=(b_{i,j,k})_{j\in A_{i},k\in m_{i,j}}$ and joint valuation profile of its fleet by $V_{i}:=(v_{i,j,k})_{j\in A_{i},k\in m_{i,j}}$. For succinct notation, we denote the joint bid and valuation profile of all FOs as $B:=(B_{i})_{i\in F}$ and $V:=(V_{i})_{i\in F}$, respectively.

### v3

@@ Abstract
The rise of advanced air mobility (AAM) is expected to become a multibillion-dollar industry in the near future. Market-based mechanisms are touted to be an integral part of AAM operations, which comprise heterogeneous operators with private valuations. In this work, we study the problem of designing a mechanism to coordinate the movement of electric vertical take-off and landing (eVTOL) aircraft, operated by multiple operators each having heterogeneous valuations associated with their fleet, between vertiports, while enforcing the arrival, departure, and parking constraints at vertiports. Particularly, we propose an incentive-compatible and individually rational vertiport reservation mechanism that maximizes a social welfare metric, which encapsulates the objective of maximizing the overall valuations of all operators while minimizing the congestion at vertiports. Additionally, we improve the computational tractability of designing the reservation mechanism by proposing a mixed binary linear programming approach that leverages the network flow structure.
@@ 1 Introduction
Despite the widespread optimism surrounding AAM, the design of regulatory policies remains an open problem. While ideas from conventional air traffic management (e.g. (Bertsimas and Patterson, 1998, 2000; Bertsimas et al., 2011; Roy and Tomlin, 2007; Odoni, 1987)) could be leveraged, they often fall short in accommodating the dynamic and adaptable nature of AAM operations (Bichler et al., 2023), resulting from on-demand requests from operators with heterogeneous private valuations (Skorup, 2019; Seuken et al., 2022). Indeed, the administrative management methods prevalent in traditional air traffic management, such as grand-fathering rights, flow management, and first-come-first-serve, prove ineffective for AAM operations (Guerreiro et al., ; Evans et al., ) as these approaches fail to elicit the heterogeneous private valuations (arising from different aircraft specifications, demand realization, etc.) different operators have on using AAM resources. Furthermore, they risk fostering inefficient and anti-competitive outcomes, as evidenced in traditional airspace operations (Dixit et al., 2023). Recognizing the need for tailored regulation, the Federal Aviation Administration (FAA) is actively developing a clean-slate congestion management framework for AAM operations to ensure efficiency, fairness, and safety (Administration, 2023).
@@
Market-based congestion management mechanisms have been proposed as potential solutions for AAM operations (Chin et al., 2023b; Qin and Balakrishnan, 2022; Wang et al., 2023; Evans et al., ; Skorup, 2019; Seuken et al., 2022).
@@
Figure 1: Schematic representation of the air traffic network with a service provider tasked with coordinating the movement of aircraft of various fleet operators between vertiports in its domain. Each vertiport has a constraint on the number of arriving aircraft, departing aircraft, and parked aircraft.
Even in conventional airspace management, market-based mechanisms are extensively studied such as (Ball et al., 2018; Basso and Zhang, 2010; Carlin and Park, 1970; Mehta and Vazirani, 2020), where both theoretical and empirical evidence show their precedence over administrative approaches (Dixit et al., 2023). However, the design of market-based mechanisms that guarantee safety, efficiency, and fairness under the heterogeneous and on-demand nature of AAM operations has remained elusive as the existing approaches concentrate heavily on tactical deconfliction (Kleinbekman et al., 2018; Bertram and Wei, ), while not accounting for efficiency, fairness and the economic incentives of operators (Chin et al., 2021; Sun et al., 2023; Chin et al., 2023a, b; Wang et al., 2023; Qin and Balakrishnan, 2022; Guerreiro et al., ; Evans et al., ).
In this paper, we introduce an auction-based mechanism for a prominent AAM scenario of vertiport reservation, where electric vertical take-off and landing (eVTOL) operators with heterogeneous private valuations need to be coordinated to use vertiports based on their realized demands. This problem is challenging for three main reasons. First, the resulting reservation must ensure efficient, fair, and safe allocation of resources. Second, the operators may misreport their private valuations and demands to gain access to more valuable airspace resources (i.e. ensuring incentive compatibility). Third, the computation of these auction mechanisms is combinatorial, as evidenced by existing air traffic flow management frameworks (Bertsimas and Patterson, 1998, 2000; Bertsimas et al., 2011) (i.e. ensuring fast computability). Thus, the main question we set out for this work is:
How to design an efficient, fair, and safe vertiport reservation mechanism for heterogeneous and on-demand nature of eVTOL operators, while ensuring incentive compatibility and faster computation?
We consider an air transportation network (ATN) managed by a service provider (SP). The SP is responsible for ensuring the efficient, safe, and fair movement of aircraft operated by various fleet operators (FOs) between vertiports (as depicted in Figure 1). The goal of the SP is to maximize a metric of social welfare that is comprised of two objectives: $(i)$ maximize the overall (weighted) valuations [^1] We allow the SP to weigh FOs differently in order to encourage new-comers in this emerging market. of all FOs, and $(ii)$ minimize excessive congestion at vertiports [^2] Note that we only consider congestion at the vertiports in this work. An extension to airborne congestion is discussed in Subsection 5.2.. Additionally, the SP must $(iii)$ enforce arrival, departure, and parking capacity constraints at vertiports, and $(iv)$ elicit truthful valuations from heterogeneous FOs in the form of bids.
@@
We propose an auction mechanism, to be used by the SP, that satisfies $(i)-(iv)$. In this mechanism, using the bids submitted by FOs, the SP allocates the resources by maximizing social welfare, subject to capacity constraints. Next, the SP charges each FO a payment based on the externality imposed by them, which is assessed by the difference in the optimal social welfare of remaining FOs when this FO is included versus when it is excluded from the auction environment. Note that this payment mechanism is inspired by the generalized Vickrey–Clarke–Groves (VCG) mechanism (Nisan et al., 2007). We theoretically study the properties of the proposed mechanism in terms of incentive compatibility, individual rationality, and social welfare maximization (cf. Theorem 3.3).
@@
There are two computational challenges associated with designing this mechanism. First, naively optimizing social welfare over the set of feasible allocations could be computationally challenging. Therefore, we frame the problem as a mixed binary linear program by constructing a network-flow graph to reduce the number of binary variables. Second, the computation of externality in the payment mechanism, which requires maximizing social welfare over the set of feasible allocations, requires characterizing the set of feasible allocations when an FO is excluded from the auction environment, which is non-trivial as the underlying resource allocation problem is an exchange problem. Therefore, we introduce the idea of pseudo-bids, where we simply set a bid of 00 to an FO while computing the optimal allocation when this FO is excluded from the auction environment.
We note two important features of the problem we study in this work. First, we focus only on strategic deconfliction where the safety is encoded in the form of minimizing congestion and ensuring capacity constraints, and not on tactical deconfliction. However, our approach can be integrated into the airborne automation workflow proposed in (Wei et al., 2023) to also account for tactical deconfliction. Second, this problem is an “exchange problem”, where some of the resources desired by any FO could be occupied by aircraft of other FOs, and a feasible allocation in this setting needs to exchange the resources between FOs while respecting capacity constraints. In constrast, the standard slot allocation problems studied in conventional air traffic literature (cf. (Dixit et al., 2023; Mehta and Vazirani, 2020; Ball et al., 2018; Rassenti et al., 1982; Pertuiset and Santos, 2014; Ball et al., 2020; Bichler et al., 2023)) are “assignment problems” where the slots need to be assigned to airlines and not exchanged between airlines.
@@
Notation: We denote the set of real numbers by $\mathbb{R}$, non-negative real numbers by $\mathbb{R}_{+}$, integers by $\mathbb{Z}$, non-negative integers by $\mathbb{Z}_{+}$, and natural numbers by $\mathbb{N}$. For $N\in\mathbb{N}$, we define $[N]:=\{1,2,...,N\}$. The indicator function is denoted as $\mathbf{1}(\cdot)$, which is 1 when ()($\cdot$)( ⋅ ) is true and 0 otherwise. When indexing a set $b=\{b_{1},b_{2},...,b_{N}\}$, we follow the standard game-theoretic notation: $b_{-i}:=\{b_{1},...,b_{i-1},b_{i+1},...,b_{N}\}$.
@@ 2.1 System Model
We consider an air transportation network (ATN), comprised of multiple vertiports, which are used by electric vertical take-off and landing (eVTOL) aircraft. We focus on a strategic deconfliction mechanism that complements the tactical deconfliction algorithms proposed in (Kleinbekman et al., 2018; Wu et al., 2022; Bertram and Wei, ; Shao et al., 2021). The scheduling mechanism proceeds over non-overlapping time slots with a receding time horizon. At the beginning of each time slot, all fleet operators (FOs) submit a menu of desired origin-destination pairs and the corresponding bids specifying how much they are willing to pay for getting scheduled. Then, the service provider (SP) will compute a feasible allocation and payment and execute them in the next time slot. The granted aircraft can now go to their desired locations. In most congested vertiports, when the parking capacity is fully utilized, any additional arrival would necessitate a simultaneous departure of an aicraft from that vertiport. Thus, this is an “exchange problem” as opposed to the “assignment problem” studied in other air traffic allocation problems (Dixit et al., 2023; Mehta and Vazirani, 2020; Ball et al., 2018; Rassenti et al., 1982; Pertuiset and Santos, 2014; Ball et al., 2020; Bichler et al., 2023).
We denote the set of vertiports by $R$, the set of FOs by $F$, and the set of eVTOL aircraft by $A$. We consider the problem for $H$ time slots.
@@
At any time $t\in[H]$, each vertiport $r\in R$ has three kinds of capacity constraints  [^3] The arrival, departure and parking capacity constraints in our model are exogeneously determined at every time step and are un-correlated between two consecutive time steps. Extending our model to account for correlations is an interesting direction of future research.: (i) arrival capacity constraints, denoted by $\textsf{arr}(r,t)\in\mathbb{Z}_{+}$, that restrict the number of eVTOLs that can land at vertiport $r$ at time $t$; (ii) departure capacity constraints, denoted by $\textsf{dep}(r,t)\in\mathbb{Z}_{+}$, that restrict the number of eVTOLs that can depart from vertiport $r$ at time $t$; (iii) parking capacity constraints, denoted by $\textsf{park}(r,t)\in\mathbb{Z}_{+}$, that restrict the number of eVTOLs that can park at vertiport $r$ at time $t$.
@@
Let $A_{i}$ be the fleet of aircraft operated by FO $i\in F$, and $A:=\{a_{i,j}|i\in F,j\in A_{i}\}$ be the set of all aircraft using the ATN. Each aircraft $a_{i,j}$ is identified by a tuple $\left(r^{\textsf{orig}}_{i,j},m_{i,j},\{t^{\textsf{dep}}_{i,j,k},t^{\textsf{arr}}_{i,j,k},v_{i,j,k},b_{i,j,k},r^{\textsf{dest}}_{i,j,k}\}_{k\in m_{i,j}}\right)$ where $(i)$ $r^{\textsf{orig}}_{i,j}\in R$ is the origin vertiport of aircraft $a_{i,j}$, $(ii)$ $m_{i,j}$ is the menu of available routes to aircraft $a_{i,j}$; $(iii)$ any route $k\in m_{i,j}$ implies that aircraft $a_{i,j}$ departs from $r^{\textsf{orig}}_{i,j}\in R$ at time $t^{\textsf{dep}}_{i,j,k}$ to arrive at $r^{\textsf{dest}}_{i,j,k}\in R$ at time $t^{\textsf{arr}}_{i,j,k}$; $(iv)$ $v_{i,j,k}$ denotes the private valuation of aircraft $a_{i,j}$ to choose the route $k\in m_{i,j}$; and $(v)$ $b_{i,j,k}\in\mathbb{R}_{+}$ is the bid submitted by FO $i$ to schedule aircraft $a_{i,j}$ on route $k\in m_{i,j}$. Note that we include the option to stay parked at the same vertiport in $m_{i,j}$, denoted by $\varnothing$, and set its departure time to 00.
Additionally, we denote the joint bid profile of all aircraft operated by FO $i\in F$ by $B_{i}:=(b_{i,j,k})_{j\in A_{i},k\in m_{i,j}}$ and joint valuation profile of its fleet by $V_{i}:=(v_{i,j,k})_{j\in A_{i},k\in m_{i,j}}$. For succinct notation, we denote the joint bid and valuation profile of all FOs as $B:=(B_{i})_{i\in F}$ and $V:=(V_{i})_{i\in F}$, respectively.